Suppose an economy has 10,000 people who are not working but looking and available for work and 90,000 people who are working. What is its unemployment rate?

Now suppose 4,000 of the people looking for work get discouraged and give up their searches. What happens to the unemployment rate? Would you interpret this as good news for the economy or bad news? Explain.


1
Expert's answer
2021-03-01T12:49:34-0500

Unemployment rate is the percentage of the total labor force that is unemployed but actively seeking employment and willing to

work.

Unemployment rate = 10,000/(90,000+10,000) = 10%

If 4,000 of the people looking for work get discouraged and give up their searches, the unemployment rate will decrease:

Unemployment rate = 6,000/(90,000+6,000)= 6.25%

It is good for the economy since less people are searching for work,

but it is bad for the economy since the total labor force decreases.